OpenStudy (anonymous):

rewrite over a common denominator: (1)/(1-sinθ)+(1)/(1+sinθ)

OpenStudy (richyw):

\[\frac{1}{1-\sin\theta}+\frac{1}{1+\sin\theta}\]

OpenStudy (anonymous):

do you know the answer?

OpenStudy (richyw):

\[=\frac{1+\sin\theta+1-\sin\theta}{\left(1-\sin\theta\right)\left( 1+\sin\theta\right)}\]\[=\frac{2}{\left(1-\sin\theta\right)\left( 1+\sin\theta\right)}\]

OpenStudy (richyw):

\[=\frac{2}{1-\sin^2\theta}\]

OpenStudy (richyw):

you can see the answer now right? it's easy from here.

OpenStudy (anonymous):

yes is it 2sec^2Theta?

OpenStudy (richyw):

yup

OpenStudy (richyw):

\[\frac{2}{\cos^2{\theta}}=2\sec^2{\theta}\]

OpenStudy (anonymous):

sinθ/cosθ+cosθ/sinθ? complete the identity

OpenStudy (richyw):

not coin it. you haven't medalled me for either question I helped you with.

OpenStudy (richyw):

not doing*

OpenStudy (anonymous):

how do i medal you im new to this?

OpenStudy (richyw):

thanks. also you are supposed to create a new question for each one! but I can help you here now anyways.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (richyw):

so multiplying to get a common denominator you have \[\frac{\sin^2x+\cos^2x}{\cos x\sin x}=\frac{1}{\cos x\sin x}=\csc x\sec x\]

OpenStudy (anonymous):

you are the best

OpenStudy (richyw):

you could also use a double angle identity to get \[2\csc{(2\theta)}\]

OpenStudy (richyw):

not sure if you have learned those or not yet. anyways I gotta run. thanks for the metals!

OpenStudy (anonymous):

oh okay thanks. I'll be needing more help in the future.

OpenStudy (anonymous):

youll be getting more from me

OpenStudy (richyw):

there are lots of people willing to help. if you see me online you can tag me using @richyw